# Asymptotic expansions for Legendre functions via differential equations having coalescing turning points

### v2

@@ Abstract
Linear second-order linear ordinary differential equations equations of the form $d^{2}w/dz^{2}=\{u^{2}f(a,z)+g(z)\}w$ are considered for large values of the real parameter $u$, $z$ being a complex variable ranging over a bounded or unbounded complex domain $Z$, and $a_{0}\leq a\leq a_{1}<\infty$. It is assumed that $f(a,z)$ and $g(z)$ are analytic in the interior of $Z$. Furthermore, $f(a,z)$ has exactly two real simple zeros in $Z$ for $a>a_{0}$ that depend continuously on $a$, and coalesce into a double zero as $a\to a_{0}$. Uniform asymptotic expansions are derived for solutions of the equation which involve parabolic cylinder functions and their derivatives, along with certain slowly varying coefficient functions. The new results involve readily computable coefficients and explicit error bounds, and are then applied to provide new asymptotic expansions for the associated Legendre functions when both the degree $\nu$ and order $\mu$ are large.
@@ 1 Introduction
For $a_{0}<a\leq a_{1}<\infty$ we assume that $f(a,z)$ has two real simple zeros at $z=z_{1}$ and $z=z_{2}$, which depend continuously on $a$, where $z_{1}<z_{2}$, and $z_{1}\to z_{2}$ as $a\to a_{0}$. We assume $f(a,z)$ and $g(z)$ are analytic in a certain domain $Z$ except possibly at the boundary of the domain. The domain $Z$ contains the two turning points, and $f(a,z)$ has no other zeros in $Z$. In addition, $f(a,z)$ is assumed to be real for real $z=x\in Z$, and $f(a,z)<0$ in $(z_{1},z_{2})$. Both $f(a,z)$ and $g(z)$ may depend on $u$ since explicit error bounds are included in the Liouville-Green (LG) expansions that we employ. Note that under the above hypotheses $f(a,z)>0$ for real $z\in Z$ such that $z<z_{1}$ and $z>z_{2}$.
@@
With this choice of $\alpha$, $\zeta$ is an analytic function of $z$ at $z=z_{j}$ ($j=1,2$). Moreover, unlike $\psi(a,z)$, one can verify that $\hat{\psi}(\alpha,\zeta)$ is also analytic at $\zeta=\pm\alpha$ (equivalently at $z=z_{1}$ and $z=z_{2}$).
@@
Note that if the turning points are bounded away from one another, expansions at each turning point could be constructed in terms of Airy functions, using the well-known theory found in [19, Chap. 11]. This results in simpler expansions, but the price paid is that the parameters are more restricted. The significance of Olver’s paper is that his approximations are uniformly valid with respect to $\alpha\in[0,A]$, and thus for the turning points bounded away from one another, as well as being allowed to coalesce into a double zero when $\alpha\to 0$ (equivalently $a\to a_{0}$). The latter situation translates into a larger permissible range for the second parameter.
In a follow up paper Olver applied the theory to Legendre functions in [16], and later to Whittaker functions [18]. We too will apply the new expansions of the present paper to Legendre functions in Section 5, which extend Olver’s one-term approximations valid for real argument to full expansions, and also that are valid for complex values of the argument. In our approach we do not consider the ODE (6), rather we use the LG transformed equation Eq. 3 only. However, the variable $\zeta$ given by Eqs. 2 and 8 still plays a central role in our analysis.
@@
We extend Olver’s general results in several ways. Our results are valid for complex values of the argument rather than just real ones, with asymptotic expansions constructed rather than just the leading term. The coefficients in our new expansions are readily computable. We mention that Olver also showed how to obtain asymptotic expansions, but these involved very complicated expressions for the coefficients, and were only shown to be valid in finite intervals, and not valid at singularities. On the other hand, our expansions are valid in unbounded domains, as well as at certain singularities of the equation. Our results are rigorously justified with explicit and relatively simple error bounds. We also asymptotically solve the connection problem, which has applications to eigenvalue problems in quantum mechanics [1].
Earlier investigations into the problem of linear second order differential equations having two turning points include those by Kazarinoff [10], Langer [11, 12], and McKelvey [14]. For other studies, see [9, 13, 15]. The results we develop here are significantly more general, and have applications to the asymptotic evaluation of various special functions, including Whittaker functions, prolate spheroidal wave functions, Bessel polynomials, and Laguerre polynomials. Physical applications include wave scattering problems for electric cylinders and prolate spheroids [10], and the eigenfunctions of an anharmonic oscillator problem in quantum mechanics [8].
The plan of this paper is as follows. In Section 2 we construct LG solutions to Eq. 1, as well as establishing important connection formulas between them. These expansions are not themselves valid at $z=z_{1}$ and $z=z_{2}$, but will be used to construct asymptotic solutions that are valid at these turning points. In Section 3 we obtain LG expansions for the parabolic cylinder functions, and these, along with the results of Section 2, are used in Section 4 to obtain our main result, namely expansions that are valid at the turning points, uniformly for $a_{0}\leq a\leq a_{1}$. In Section 5 apply our new theory to obtain uniform asymptotic expansions for the associated Legendre functions when both the degree $\nu$ and the order $\mu$ are large. Finally in Section 6 we illustrate the accuracy of our uniform expansions with some numerical results from Section 5 for two cases, where the turning points in question are both close and not close.
@@ 2 LG expansions for the general equation
Let points $z=x_{j}$ ($j=1,2,3,4$) in the $z$ plane correspond to $\zeta=i\infty$, +$\infty$+ ∞, $-i\infty$, and $\infty$- ∞, respectively. Since $\zeta$ is unbounded at these points, the corresponding $z$ points must also be unbounded, or at a finite point that is a singularity of (1). We assume that all four are boundary points of $Z$.
@@
For $u>0$ sufficiently large, let $k_{n,j}(u,a)$ ($j=1,2,3$) be appropriately chosen continuous functions of $a$ for $a_{0}\leq a\leq a_{1}$. Then we apply [4, Thm. 1.1] to obtain the LG expansions to the solutions of (1) that are recessive at $z=x_{j}$ ($j=1,2,3$), given by
(18) $w_{n,j}(u,a,z)=\frac{k_{n,j}(u,a)W_{n,j}(u,a,z)}{f^{1/4}(a,z)}\quad(j=1,2,3),$
where, for an arbitrary positive integer $n$,
@@
(29) $w_{n,4}(u,a,z)=\frac{k_{n,4}^{\pm}(u,a)W_{n,4}^{\pm}(u,a,z)}{f^{1/4}(a,z)},$
Let us now examine the connection formulas between the four fundamental solutions. From linear second-order differential equation theory, we know any one of them can be expressed as a linear combination of two of the others. With this in mind, let the constants $k_{n,j}(u,a)$ ($j=1,2,3$) of Eqs. 18 and 29 be scaled relative to each other so that the linear relationship between $w_{n,j}(u,a,z)$ ($j=1,2,3$) is given by
(30) $w_{n,2}(u,a,z)=w_{n,1}(u,a,z)+w_{n,3}(u,a,z).$
Given these values, we then express the linear relation between $w_{n,j}(u,a,z)$ ($j=1,3,4$) in the form
(31) $w_{n,4}(u,a,z)=e^{\lambda\pi i}w_{n,1}(u,a,z)+e^{-\lambda\pi i}w_{n,3}(u,a,z),$
@@
(33) $2i\sin(\lambda\pi)w_{n,3}(u,a,z)=e^{\lambda\pi i}w_{n,2}(u,a,z)-w_{n,4}(u,a,z).$
As $z\to x_{1}$ we have $w_{n,1}(u,a,z)\to 0$ and $w_{n,j}(u,a,z)$ ($j=2,4$) unbounded. Now, the LG expansions Eqs. 19 and 28 are both asymptotically valid at $z=x_{1}$, provided that in the latter upper signs are taken. Thus, with these expansions, along with (32), we deduce that
@@
(38) $e^{2\lambda\pi i}=-\lim_{z\to x_{4}}\frac{w_{n,3}(u,a,z)}{w_{n,1}(u,a,z)}.$
Then in this we use the LG expansions (19) and taking $z\to x_{4}\pm i0$ ($\xi\to\infty\mp\frac{1}{2}\alpha^{2}\pi i$) for $w_{n,1}(u,a,z)$ and $w_{n,3}(u,a,z)$, respectively. Then on referring to (36), for any positive integer $n$, we arrive at
(39) $e^{2\lambda\pi i}=-\exp\left\{u\alpha^{2}\pi i+\sum\limits_{s=1}^{n-1}{\frac{\hat{E}_{s}\left(a,x_{4}-i0\right)-\hat{E}_{s}\left(a,x_{4}+i0\right)}{u^{s}}}\right\}\\ \times\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty).$
Now from Eqs. 4, 22, 23, and 24 one can show by induction that the even coefficients $\hat{E}_{2s}(a,z)$ ($s=1,2,3,\dots$) are meromorphic at the turning points $z=z_{j}$ ($j=1,2$), and so $\hat{E}_{2s}(a,x_{4}+i0)=\hat{E}_{2s}(a,x_{4}-i0)$. Thus (39) yields Eq. 37 (modulo an arbitrary integer, which we take to be zero)
@@
For arbitrary $\delta>0$ assume that $a_{0}+\delta\leq a\leq a_{1}$. Then
(40) $\frac{k_{n,1}(u,a)}{k_{n,2}(u,a)}\sim i\exp\left\{\sum_{s=0}^{\infty}\frac{\kappa_{2s+1}(a)}{u^{2s+1}}\right\}\quad(u\to\infty),$
for certain constants $\kappa_{2s+1}(a)$ which depend on the choice of integration constants of the odd terms in (21).
@@
The assumption $a_{0}+\delta\leq a\leq a_{1}$ implies that $z_{1}$ is bounded away from $z_{2}$. Then we can follow the steps of the proof of [7, Lemma 2.3], with the turning point $z_{2}$ of the present paper playing the role of $z_{0}$ of that paper, and $w_{n,j}(u,a,z)$ ($j=1,2,3$) here replacing $W_{k}(u,z)$ ($k=1,0,-1$) respectively. This establishes that (40) is true.
Remark 2.5.
The constants on the RHS of (40) are in effect zero in the result of [7, Lemma 2.3], since in that paper the integration constants for the odd LG coefficients ([7, Eq. (1.14)]) are chosen so that they are $(z-z_{0})^{1/2}$ times a meromorphic function at the turning point $z_{0}$, which is not generally the case for the odd LG coefficients given by (21) at our turning point $z_{2}$
Remark 2.6.
Under certain conditions on the LHS of Eq. 40 and on the behavior of the zeros $z_{j}$ ($j=1,2$) as functions of $a$ it is possible to extend Eq. 40 to $a_{0}\leq a\leq a_{1}$, but the proof would be quite involved, and so we do not pursue this here. Instead, in most applications, one can rely on explicit or known asymptotic representations of the constants $k_{n,j}(u,a)$ ($j=1,2$) to establish this extension if it holds.

### v3

@@ Abstract
Linear second-order linear ordinary differential equations equations of the form $d^{2}w/dz^{2}=\{u^{2}f(a,z)+g(z)\}w$ are considered for large values of the real parameter $u$, $z$ being a complex variable ranging over a bounded or unbounded complex domain $Z$, and $a_{0}\leq a\leq a_{1}<\infty$. It is assumed that $f(a,z)$ and $g(z)$ are analytic in the interior of $Z$. Furthermore, $f(a,z)$ has exactly two real simple zeros in $Z$ for $a>a_{0}$ that depend continuously on $a$, and coalesce into a double zero as $a\to a_{0}$. Uniform asymptotic expansions are derived for solutions of the equation which involve parabolic cylinder functions and their derivatives, along with certain slowly-varying coefficient functions. The new results involve readily computable coefficients and explicit error bounds, and are then applied to provide new asymptotic expansions for the associated Legendre functions when both the degree $\nu$ and order $\mu$ are large.
@@ 1 Introduction
For $a_{0}<a\leq a_{1}<\infty$ we assume that $f(a,z)$ has two real simple zeros at $z=z_{1}$ and $z=z_{2}$, which depend continuously on $a$, where $z_{1}<z_{2}$, and $z_{1}\to z_{2}$ as $a\to a_{0}$. We assume $f(a,z)$ and $g(z)$ are analytic in a certain domain $Z$ except possibly at the boundary of the domain. The domain $Z$ contains the two turning points, and $f(a,z)$ does not have other zeros in $Z$. In addition, $f(a,z)$ is assumed to be real for real $z=x\in Z$, and $f(a,z)<0$ in $(z_{1},z_{2})$. Both $f(a,z)$ and $g(z)$ may depend on $u$ since explicit error bounds are included in the Liouville-Green (LG) expansions that we employ. Note that under the above hypotheses $f(a,z)>0$ for real $z\in Z$ such that $z<z_{1}$ and $z>z_{2}$.
@@
With this choice of $\alpha$, $\zeta$ is an analytic function of $z$ at $z=z_{j}$ ($j=1,2$). Moreover, unlike $\psi(a,z)$, one can verify that $\hat{\psi}(\alpha,\zeta)$ is also analytic at $\zeta=\pm\alpha$ (equivalently at $z=z_{1}$ and $z=z_{2}$). Note that $\alpha\to 0$ as $a\to a_{0}$, and so we let $0\leq\alpha\leq\alpha_{1}$ correspond to $a_{0}\leq a\leq a_{1}$.
@@
Note that if the turning points are bounded away from one another, expansions at each turning point could be constructed in terms of Airy functions, using the well-known theory found in [19, Chap. 11]. This results in simpler expansions, but the price paid is that the parameters are more restricted. The significance of Olver’s paper is that his approximations are uniformly valid with respect to $\alpha\in[0,\alpha_{1}]$, and thus for the turning points bounded away from one another, as well as being allowed to coalesce into a double zero when $\alpha\to 0$ (equivalently $a\to a_{0}$). The latter situation translates into a larger permissible range for the second parameter.
In a follow-up paper Olver applied the theory to Legendre functions in [16], and later to Whittaker functions [18]. We too will apply the new expansions of the present paper to Legendre functions in Section 5, which extend Olver’s one-term approximations valid for real argument to full expansions, and also that are valid for complex values of the argument. In our approach we do not consider the ODE (6), rather we use the LG transformed equation Eq. 3 only. However, the variable $\zeta$ given by Eqs. 2 and 8 still plays a central role in our analysis.
@@
We extend Olver’s general results in several ways. Our results are valid for complex values of the argument rather than just real ones, with asymptotic expansions constructed rather than just the leading term. The coefficients in our new expansions are readily computable. We mention that Olver also showed how to obtain asymptotic expansions, but these involved very complicated expressions for the coefficients, were only shown to be valid in finite intervals, and are not valid at singularities. On the other hand, our expansions are valid in unbounded domains, as well as at certain singularities of the equation. Our results are rigorously justified with explicit and relatively simple error bounds. We also asymptotically solve the connection problem, which has applications to eigenvalue problems in quantum mechanics [1].
Earlier investigations into the problem of linear second-order differential equations having two turning points include those by Kazarinoff [10], Langer [11, 12], and McKelvey [14]. For other studies, see [9, 13, 15]. The results we develop here are significantly more general, and have applications to the asymptotic evaluation of various special functions, including Whittaker functions, prolate spheroidal wave functions, Bessel polynomials, and Laguerre polynomials. Physical applications include wave scattering problems for electric cylinders and prolate spheroids [10], and the eigenfunctions of an anharmonic oscillator problem in quantum mechanics [8].
The plan of this paper is as follows. In Section 2 we construct LG solutions for Eq. 1, as well as establishing important connection formulas between them. These expansions are not themselves valid at $z=z_{1}$ and $z=z_{2}$, but will be used to construct asymptotic solutions that are valid at these turning points. In Section 3 we obtain LG expansions for the parabolic cylinder functions, and these, together with the results of Section 2, are used in Section 4 to obtain our main result, namely expansions that are valid at the turning points, uniformly for $a_{0}\leq a\leq a_{1}$. In Section 5 apply our new theory to obtain uniform asymptotic expansions for the associated Legendre functions when both the degree $\nu$ and the order $\mu$ are large. Finally in Section 6 we illustrate the accuracy of our uniform expansions with some numerical results from Section 5 for two cases, where the turning points in question are both close and not close.
@@ 2 LG expansions for the general equation
Let points $z=x_{j}$ ($j=1,2,3,4$) in the $z$ plane correspond to $\zeta=i\infty$ ($\Re(\xi)=-\infty$), $\zeta=+\infty$ ($\Re(\xi)=+\infty$), $\zeta=-i\infty$ ($\Re(\xi)=-\infty$), and $\zeta=-\infty$ ($\Re(\xi)=+\infty$), respectively. Since $\zeta$ is unbounded at these points, the corresponding $z$ points must also be unbounded, or at a finite point that is a singularity of (1). We assume that all four are boundary points of $Z$.
@@
We shall call such paths $\mathcal{L}_{j}$ progressive.
For each positive integer $n$, and $u>0$ sufficiently large, let $k_{n,j}(u,a)$ ($j=1,2,3$) be appropriately chosen continuous functions of $a$ for $a_{0}\leq a\leq a_{1}$. Then we apply [4, Thm. 1.1] to obtain the LG expansions to the solutions of (1) that are recessive at $z=x_{j}$ ($j=1,2,3$), given by
(18) $w_{j}(u,a,z)=\frac{k_{n,j}(u,a)W_{n,j}(u,a,z)}{f^{1/4}(a,z)}\quad(j=1,2,3),$
where
@@
The coefficients $k_{n,j}(u,a)$ are chosen relative to certain connection formulas which we discuss later, and are such that each $w_{j}(u,a,z)$ ($j=1,2,3$) is independent of $n$ (and likewise for $w_{4}(u,a,z)$ given by Eq. 29 below).
@@
(29) $w_{4}(u,a,z)=\frac{k_{n,4}^{\pm}(u,a)W_{n,4}^{\pm}(u,a,z)}{f^{1/4}(a,z)},$
Let us now examine the connection formulas between the four fundamental solutions. From linear second-order differential equation theory, we know that any one of them can be expressed as a linear combination of two of the others that are linearly independent. With this in mind, let the constants $k_{n,j}(u,a)$ ($j=1,2,3$) of Eqs. 18 and 29 be scaled relative to each other so that the linear relationship between $w_{j}(u,a,z)$ ($j=1,2,3$) is given by
(30) $w_{2}(u,a,z)=w_{1}(u,a,z)+w_{3}(u,a,z),$
noting that $w_{1}(u,a,z)$ and $w_{3}(u,a,z)$ are linearly independent for our parameter range.
Given these values, we then express the linear relation between $w_{j}(u,a,z)$ ($j=1,3,4$) in the form
(31) $w_{4}(u,a,z)=e^{\lambda\pi i}w_{1}(u,a,z)+e^{-\lambda\pi i}w_{3}(u,a,z),$
@@
(33) $2i\sin(\lambda\pi)w_{3}(u,a,z)=e^{\lambda\pi i}w_{2}(u,a,z)-w_{4}(u,a,z).$
As $z\to x_{1}$ we have $w_{1}(u,a,z)\to 0$ and $w_{j}(u,a,z)$ ($j=2,4$) unbounded. Now, the LG expansions Eqs. 19 and 28 are both asymptotically valid at $z=x_{1}$, provided that in the latter upper signs are taken. Thus, with these expansions, along with (32), we deduce that
@@
(38) $e^{2\lambda\pi i}=-\lim_{z\to x_{4}}\frac{w_{3}(u,a,z)}{w_{1}(u,a,z)}.$
In this we use the LG expansions (19) and take $z\to x_{4}\pm i0$ ($\xi\to\infty\mp\frac{1}{2}\alpha^{2}\pi i$) for $w_{1}(u,a,z)$ and $w_{3}(u,a,z)$, respectively. Then on referring to (36), for any positive integer $n$, we arrive at
(39) $e^{2\lambda\pi i}=-\exp\left\{u\alpha^{2}\pi i+\sum\limits_{s=1}^{n-1}{\frac{\hat{E}_{s}\left(a,x_{4}-i0\right)-\hat{E}_{s}\left(a,x_{4}+i0\right)}{u^{s}}}\right\}\\ \times\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty),$
on noting that
(40) $\frac{\lim_{z\to x_{4}-i0}f^{-1/4}(u,z)}{\lim_{z\to x_{4}+i0}f^{-1/4}(u,z)}=-1,$
which follows from $f^{-1/4}(u,z)=\{(z-z_{1})(z-z_{2})\}^{-1/4}p(a,z)$, where $p(a,z)$ is analytic in $Z$. Now, from Eqs. 4, 22, 23, and 24, one can show by induction that the even coefficients $\hat{E}_{2s}(a,z)$ ($s=1,2,3,\dots$) are meromorphic at the turning points $z=z_{j}$ ($j=1,2$), and so $\hat{E}_{2s}(a,x_{4}+i0)=\hat{E}_{2s}(a,x_{4}-i0)$. Thus (39) yields Eq. 37 (modulo an arbitrary integer, which we take to be zero)
@@
For arbitrary $\delta>0$ assume that $a_{0}+\delta\leq a\leq a_{1}$. Then for any positive integer $n$
(41) $\frac{k_{n,1}(u,a)}{k_{n,2}(u,a)}=i\exp\left\{\sum_{s=0}^{n-1}\frac{\kappa_{2s+1}(a)}{u^{2s+1}}\right\}\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty),$
for certain constants $\kappa_{2s+1}(a)$ which depend on the odd coefficients in (21).
@@
We use (30), and in this relation let $z\to x_{3}$, in which case $w_{3}(u,a,z)\to 0$, while the other two functions grow exponentially, but whose ratio is bounded in this limit. We use the LG expansions Eqs. 18, 19, and 20 for $w_{1}(u,a,z)$ and $w_{2}(u,a,z)$ for an arbitrary positive integer $n$. To this end, observe from Eq. 9 that $\xi\to-\infty+\frac{1}{4}\alpha^{2}\pi i$ as $z\to x_{3}$ ($\zeta\to-i\infty$). This can be used for the LG expansion of $w_{2}(u,a,z)$, since $x_{3}$ lies in the LG region of validity, $Z_{2}$, for $W_{n,2}(u,z)$.
The determination of the corresponding LG limit as $z\to x_{3}$ for $w_{1}(u,a,z)$ is not quite as straightforward, since the variable must be accessed across the cut $[z_{1},z_{2}]$ in the $z$ plane, as this is the only way to enter the lower half-plane along a progressive path $\mathcal{L}_{1}$ used in the error bound. This path can then be chosen to satisfy the condition that $\Re(u\xi(v))$ be monotonic for points $v$ moving along the path from the end points $x_{1}$ to $x_{3}$. In addition, the path must be bounded away from both turning points as it passes through $[z_{1},z_{2}]$, and that is why we require them to be bounded away from each other, given by the condition $a_{0}<a_{0}+\delta\leq a$.
Consider then $z\to x_{3}$ from crossing the part of the cut $[z_{1},z_{2}]$ from above, and label this $z=x_{3}^{-}$. Then from Eq. 9 we find $\xi=\to+\infty-\frac{1}{4}\alpha^{2}\pi i$ as $z\to x_{3}^{-}$ on this sheet. Thus, on using the fact that $w_{3}(u,a,z)/w_{2}(u,a,z)\to 0$ as $z\to x_{3}$, we have from Eqs. 18 and 30
(42) $1=\lim_{z\to x_{3}}\frac{w_{1}(u,a,z)}{w_{2}(u,a,z)}=\frac{k_{n,1}(u,a)\lim_{z\to x_{3}^{-}}f^{-1/4}(u,z)W_{n,1}(u,a,z)}{k_{n,2}(u,a)\lim_{z\to x_{3}}f^{-1/4}(u,z)W_{n,2}(u,a,z)}.$
Next, similarly to Eq. 40, we have
(43) $\frac{\lim_{z\to x_{3}^{-}}f^{-1/4}(u,z)}{\lim_{z\to x_{3}}f^{-1/4}(u,z)}=-i.$
Further, since $\xi=\to\pm\infty\mp\frac{1}{4}\alpha^{2}\pi i$ as $z\to x_{3}^{-}$ and $z\to x_{3}$, respectively, from Eqs. 19 and 20 it is seen that
(44) $\frac{\lim_{z\to x_{3}^{-}}W_{n,1}(u,a,z)}{\lim_{z\to x_{3}}W_{n,2}(u,a,z)}=\exp\left\{\sum\limits_{s=1}^{n-1}{\frac{\hat{E}_{s}(a,x_{3}^{-})+(-1)^{s-1}\hat{E}_{s}(a,x_{3})}{u^{s}}}\right\}\\ \times\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty).$
Since the even coefficients $\hat{E}_{2s}(a,z)$ ($s=1,2,3,\dots$) are meromorphic at the turning points $z=z_{j}$ ($j=1,2$), and therefore single-valued at $z=x_{3}$, it follows that $\hat{E}_{2s}(a,x_{3}^{-})-\hat{E}_{2s}(a,x_{3})=0$. Thus, Eqs. 42, 43, and 44 yields Eq. 41, where $\kappa_{2s+1}(a)=-\{\hat{E}_{2s+1}(a,x_{3}^{-})+\hat{E}_{2s+1}(a,x_{3})\}$.